OpenStudy (dallaslifebaby01):

Solve 4x – 5 = 11. Show each step of your work.

OpenStudy (cggurumanjunath):

add 5 on both sides

OpenStudy (calculusxy):

\(\large 4x - 5 = 11\) Add \(\large 5\) to both sides to isolate x. \(\large 4x - 5 + 5 = 11 + 5\) Now you have: \(\large 4x = 16\) Use the division property of equality by dividing 4 both sides. \(\large \frac{4x}{4} = \frac{16}{4}\) Then you have: \(\large x = 4\)

OpenStudy (cggurumanjunath):

4x-5+5=11+5

OpenStudy (dallaslifebaby01):

okay thanks i wasn't sure if i did it right! i have 2 more questions could you help me with those to?

OpenStudy (calculusxy):

Can try ..

OpenStudy (cggurumanjunath):

we add 5 to isolate the variable.

OpenStudy (dallaslifebaby01):

thanks

OpenStudy (dallaslifebaby01):

one second

OpenStudy (dallaslifebaby01):

Solve 5x + 4 = 3x – 8. Show each step of your work.

OpenStudy (cggurumanjunath):

is that step clear ?

OpenStudy (dallaslifebaby01):

i understand the first one and i had previously worked it out but i wasnt sure if i did it correctly

OpenStudy (calculusxy):

It's basically the same thing but this time you need to combine the like terms, meaning the \(x\) first. \(\large 5x + 4 = 3x - 8\) Put the coefficients with the \(\large x\) together. \(\large 5x - 3x + 4 = -8\) Now you have: \(\large 2x + 4 = -8\) Do the same things as above now: \(\large 2x + 4 - 4 = -8 - 4\) Penultimately, you have: \(\large 2x = -12\) Divide both sides by 2 to get both: \(\large x = -6\) You can plug in \(\large -6\) for \(\large x\) in the equation to see whether it is correct or not.

OpenStudy (dallaslifebaby01):

okay the last one is Solve 7(x + 8) + 2(x – 4) = 12. Show each step of your work.

OpenStudy (calculusxy):

Exactly the same thing now with distributive property: \(7x + 56 + 2x - 8 = 12\) I think you should be able to solve it from here. Use previous posts as references if needed.

OpenStudy (dallaslifebaby01):

okay thanks

OpenStudy (calculusxy):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @calculusxy It's basically the same thing but this time you need to combine the like terms, meaning the \(x\) first. \(\large 5x + 4 = 3x - 8\) Put the coefficients with the \(\large x\) together. \(\large 5x - 3x + 4 = -8\) Now you have: \(\large 2x + 4 = -8\) Do the same things as above now: \(\large 2x + 4 - 4 = -8 - 4\) Penultimately, you have: \(\large 2x = -12\) Divide both sides by 2 to get both: \(\large x = -6\) You can plug in \(\large -6\) for \(\large x\) in the equation to see whether it is correct or not. \(\color{#0cbb34}{\text{End of Quote}}\) Use this one especially...

OpenStudy (calculusxy):

@DallasLifeBaby01 You're welcome :)